# Responsible AI Question Bank: A Comprehensive Tool for AI Risk Assessment

### v1

@@ 1 Introduction
A recent report found that while many companies view AI as a promising technology and actively pursue AI opportunities, only 10% of those surveyed have publicly announced their Responsible AI (RAI) policies  [^1] https://www.csiro.au/-/media/D61/Responsible-AI/Alphinity/Responsible-AI-and-ESG.pdf. This suggests that many companies are still concerned about their RAI maturity level. Moreover, there have been numerous incidents involving AI  [^2] https://incidentdatabase.ai/. AI incidents are not limited to specific industries like finance and healthcare; they have emerged across all sectors, raising various concerns in areas such as privacy, bias, and safety.
@@
There are a number of industrial RAI frameworks or AI governance frameworks to support the implementation of RAI principles in practice. Most of them are principle-driven and process-centred frameworks which provide a set of guidance for evaluating RAI practices and mitigating potential risks associated with AI systems. In the US, the NIST AI Risk Management Framework  [^3] https://www.nist.gov/itl/ai-risk-management-framework provides four management functions such as MAP, GOVERN, MEASURE and METRIC to manage AI risks . Recently, NIST has also released a new draft framework for GenAI  [^4] https://airc.nist.gov/docs/NIST.AI.600-1.GenAI-Profile.ipd.pdf, which encompasses 12 potential risks of generative AI (GenAI) and provides comprehensive RAI practices to address those risks. In Europe, the EU Assessment List for Trust AI Framework  [^5] https://futurium.ec.europa.eu/en/european-ai-alliance/pages/welcome-altai-portal has been used as AI ethics guidelines, which aligns with seven RAI principles (e.g., privacy, transparency, accountability) . In academia, researchers have increasingly focused on RAI practices and risk assessment for AI.
@@
However, there is a lack of a comprehensive and well-structured AI risk assessment. Especially, there is absence or a lack of a RAI assessment toolkit that incorporates interconnected and layered risk questions and connectivity to other standards and frameworks [3]. Existing RAI toolkits often focus on specific aspects, overlooking the holistic nature of RAI principles. For instance, Liao et al. [4, 5] introduced a question bank specifically focused on transparency and explainability, neglecting other crucial principles. Additionally, the US Department of Energy’s AI Risk Management Playbook  [^6] https://www.energy.gov/ai/doe-ai-risk-management-playbook-airmp provides extensive guidance on identifying AI risks and recommending mitigations, but it lacks integration with diverse standards, frameworks, and related resources.
@@
To address current challenges, we have developed a novel RAI question bank since 2022. The initial version provided a general framework suitable for broad application, though it lacked the specific focus for more specialized use cases [6]. A key advancement is the integration of a systematic approach that links lower-level risk questions to higher-level ones and related themes. This approach prevents siloed and fragmented assessments, ensuring a cohesive evaluation process across different levels of the organization. Additionally, our question bank is designed to accommodate different levels of stakeholders (e.g., executives, senior managers, developers) and is mapped to the system lifecycle stages. This design allows users to focus on their specific contexts and interests while maintaining a comprehensive system perspective, thus enhancing both the relevance and effectiveness of the risk assessments. Recent enhancements have significantly increased its flexibility, enabling support for a wide range of use case scenarios. These include compliance checks with various policies, such as the EU AI Act and ISO/IEC AI management standards, as well as risk assessments for AI agents and foundation models.
@@
While the RAI Question Bank primarily uses yes/no questions, which may lead to a checkbox mentality (i.e., merely the existence of very low-quality artifacts/processes) if not carefully managed, we acknowledge the need for quality metrics and evidence-based answers. This means that the questions need to have quality aspects and answers are not just yes/no but require evidence, and lower-level answers collectively provide a crude metric for higher-level questions. Although quality metrics and measurement are not the primary focus of this study, we have examined these considerations through a case study (Section 5.2) where we tested and refined the approach to ensure more meaningful and reliable assessments.
This study aims to introduce our Responsible AI Question Bank, serving as a fundamental resource and tool for AI risk assessment. We present two case studies to illustrate the practical application of the RAI Question Bank and discuss stakeholder feedback and implications for real-world use. We also discuss how our question bank can be used to comply with regulations and build trust in the age of AI agents like Auto-GPT  [^7] https://github.com/Significant-Gravitas/Auto-GPT. AI agents, especially, foundation model-based agents go beyond simple chatbots, autonomously achieving human goals. This expanded capability and autonomy require stricter standards for responsible AI development and accountability for their actions [7].
@@
The rest of the paper comprises as follows. In Section 2, we provide background information and literature review results. Section 3 describes the overall design and development approach of the RAI Question Bank. We then introduce the overview and details of the RAI Question Bank in the next section (Section 4). Section 5 includes two case studies such as AI risk assessment for AI projects and ESG-AI investor framework. Section 6 discusses the practical implications of this study, including the potential use of the RAI Question Bank for compliance checking and risk assessment of AI agents. We conclude this study in Section 7.
2 Background and Literature Review
@@
    NIST AI Risk Management Framework (AI RMF)3: The NIST AI RMF offers a flexible and voluntary framework applicable across various sectors and use cases. It enhances AI system trustworthiness by focusing on four main functions: Govern, Map, Measure, and Manage. Developed through an inclusive and transparent process, the framework supports existing AI risk management initiatives and adapts to the evolving AI landscape, enabling organizations to implement it to different extents for societal benefit while minimizing risks.
  2. 2.
@@
    Microsoft Responsible AI Impact Assessment Template [^8] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2022/06/Microsoft-RAI-Impact-Assessment-Template.pdf and Guide [^9] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2022/06/Microsoft-RAI-Impact-Assessment-Guide.pdf: Microsoft’s resources support its Responsible AI Standard [^10] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2022/06/Microsoft-Responsible-AI-Standard-v2-General-Requirements-3.pdf. They enable AI development teams to explore and understand the impacts of their AI systems on stakeholders, intended benefits, and potential harms from the earliest design stages.
  3. 3.
@@
    EU Assessment List for Trustworthy Artificial Intelligence (ALTAI) [^11] https://digital-strategy.ec.europa.eu/en/library/assessment-list-trustworthy-artificial-intelligence-altai-self-assessment: ALTAI helps organizations self-assess the trustworthiness of their AI systems. This checklist translates key requirements from the Ethics Guidelines for Trustworthy AI [^12] https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai into actionable steps, ensuring alignment with principles to mitigate risks and promote responsible AI development and deployment.
  4. 4.
@@
    Canada Algorithmic Impact Assessment (AIA) [^13] https://www.canada.ca/en/government/system/digital-government/digital-government-innovations/responsible-use-ai/algorithmic-impact-assessment.html: This tool helps organizations evaluate the potential impacts of AI systems, especially in public services. It includes a detailed questionnaire to identify and mitigate risks associated with automated decision systems, focusing on bias, privacy, and transparency, and ensures compliance with ethical and legal standards.
  5. 5.
@@
    Australia NSW AI Assurance Framework [^14] https://www.digital.nsw.gov.au/sites/default/files/2022-09/nsw-government-assurance-framework.pdf: This framework ensures the safe, ethical, and responsible use of AI technologies in government services. It includes AI governance, risk management, and mitigation strategies, emphasizing stakeholder engagement and continuous monitoring to maintain compliance with ethical standards. It also addresses emerging AI risks, such as those posed by generative AI, ensuring AI projects meet community expectations and uphold public trust.
AI risk assessments should account for the evolving landscape of AI regulations and standards. EU AI Act  [^15] https://www.europarl.europa.eu/doceo/document/TA-9-2024-0138-FNL-COR01_EN.pdf and ISOIEC 42001:2023  [^16] https://www.iso.org/standard/81230.html are emerging AI regulation and standard to be considered for AI risk assessment.
The EU AI Act establishes a comprehensive regulatory framework for AI within the European Union. It classifies AI systems based on their risk levels: unacceptable, high, limited, and minimal risk. Unacceptable risk AI applications are banned, while high-risk applications must meet stringent requirements for transparency, quality, and security, and undergo conformity assessments. Additionally, it establishes the European Artificial Intelligence Board to ensure compliance and promote cooperation among member states.
@@
ISO/IEC 42001:2023 specifies requirements for establishing, implementing, maintaining, and continually improving an AI Management System (AIMS) within organizations. It is designed for entities providing or utilizing AI-based products or services, ensuring responsible development and use of AI. The standard focuses on quality management and reliability, guiding organizations in managing the issues and risks associated with AI technologies. This systematic approach facilitates consistency with other management system standards related to quality, safety, security, and privacy.
@@
Recent efforts to improve the safety of advanced AI include several notable works. Partnership on AI released guidance for the safe deployment of foundation models, targeting model developers [^17] https://partnershiponai.org/modeldeployment/. The AI Verify Foundation and Singapore’s Infocomm Media Development Authority published the Model AI Governance Framework for Generative AI, providing high-level guidance on the design, development, and deployment of generative AI models [^18] https://aiverifyfoundation.sg/wp-content/uploads/2024/05/Model-AI-Governance-Framework-for-Generative-AI-May-2024-1-1.pdf. On April 29, 2024, NIST released a draft of the AI RMF Generative AI Profile, guiding organizations in managing risks unique to generative AI [^19] https://airc.nist.gov/docs/NIST.AI.600-1.GenAI-Profile.ipd.pdf. The EU AI Act also covers general-purpose AI like generative models, imposing specific transparency and evaluation requirements. The UK AI Safety Institute has published several resources, including Inspect [^20] https://github.com/UKGovernmentBEIS/inspect_ai, a platform for evaluating LLMs, and is actively developing infrastructure to assess the safety and societal impacts of advanced AI. Recently, the AU NSW Government updated its Artificial Intelligence Assessment Framework [^21] https://www.digital.nsw.gov.au/policy/artificial-intelligence/nsw-artificial-intelligence-assessment-framework. The revised framework, formerly known as the NSW Artificial Intelligence Assurance Framework, now categorizes generative AI-based solutions as "elevated risk," requiring additional and stricter assessments.
@@ 3 Methodology
This section introduces the methodology of this study. As shown in Figure 1, this study was conducted in five dedicated phases from 2022 to 2024.
First, a systematic mapping study was performed to understand the state-of-the-art in AI risk assessment and select reference frameworks for developing the RAI question bank  [3]. We selected five frameworks from the results, scrutinized, and extracted all included AI risk questions. The collected questions were synthesized and used to develop a comprehensive and holistic question bank for AI risk assessment [6]. We then conducted a case study with eight AI development projects. Some selected questions were tailored for each project and used for interviews with team members. We gained insights and feedback from the participating project teams. Based on the feedback, we improved the initial version of the question bank. As new AI regulations and standards are widely adopted, we incorporated the EU AI Act and ISO/IEC AI management standard to support organizations in achieving compliance. Finally, we evaluated the feasibility and applicability of the question bank through case studies. These studies included the eight AI projects from the initial case studies and the ESG-AI framework development project, which involved engagement with 28 companies that have adopted AI technology in their business workflows.
Figure 1: Research steps overview.
@@ 3.2 Question Bank Development
Based on the previous phase, five frameworks out of the 16 frameworks was chosen for further analysis and synthesis. The frameworks were selected according to i) global recognition, ii) inclusion of risk assessment questions, and, iii) representation of different regions and industry leaders. The selected frameworks include: US NIST AI Risk Management Framework3, EU Trustworthy AI Assessment List12, Canada Algorithmic Impact Assessment13, Australia NSW AI Assurance Framework14, and Microsoft Responsible AI Impact Assessment Guide9. By investigating these frameworks, we can comprehend and synthesize risk assessment solutions throughout various geopolitical contexts.
We began by extracting 382 questions for risk assessment, which were subsequently refined through an iterative process to consolidate questions on similar topics. This refinement resulted in a condensed question bank of 293 questions. We utilized Australia’s AI ethics principles  [^22] https://www.industry.gov.au/publications/australias-artificial-intelligence-ethics-framework/australias-ai-ethics-principles to categorize these questions. We structured the hierarchy of our question bank using a decision tree [20], organizing questions based on their levels (level 1-3) and sequences to ensure a coherent and comprehensive framework. To address potential gaps in technical/practical detail at lower level questions, we recognized that some lower-level (level 2 and 3) questions can be inferred from broader high-level frameworks. We made inferences to generate or extract these detailed questions where necessary. Concept mapping [21] was employed to identify common risk themes, allowing us to further analyze and group the questions.
@@ 3.3 Case Study
The initial case study (Phase 3) with eight AI projects was conducted in 2023. We analyzed these projects to understand their contexts and selected 3 to 5 questions from our question bank for each AI ethics principle to prepare for interviews. In parallel, we prepared a dedicated risk register template, including risk ID, category, title, description, causes, and selected interview questions derived from the question bank. For each project, we scheduled a 1.5-hour interview. We synthesized and analyzed the interview notes to identify and assess the AI ethics risks the project teams might encounter. We also asked the interviewees for feedback on the question bank for further refinement. We conducted two rounds of interviews in 2023 and consolidated the feedback to improve our question bank.
The second case study (Phase 5) was conducted with the ESG-AI investor framework development project, following the question bank improvement (Phase 3), based on the initial case study results and changes in the external AI regulation landscape. This case study engaged 28 companies using and developing AI. Considering the companies’ characteristics and contexts, we selected and customized 5-10 questions for the interviews. In the ESG-AI framework development, 42 questions from the RAI Question Bank were selected to contribute to a deep dive assessment of companies. During the project, we received user feedback, improved the RAI Question Bank accordingly, and discussed the next steps.
Section 5 provides details about the case studies, user feedback, and improvements.
@@ 4.1 Overview
This section introduces the RAI Question Bank which is a set of risk questions designed to support the ethical development of AI systems. It addresses the needs of various users, including C-level executives, senior managers, developers, and others.
@@
Figure 2 presents the overall architecture of the RAI Question Bank. It is based on AI ethics principles which serve as an overarching category for the question bank. Specifically, we have selected Australia’s AI ethics principles22 to comprehensively cover responsible AI concerns. The principles closely align with similar principles [22, 23] from around the world. The principles comprise eight principles such as Human, Societal and Environmental Wellbeing, Human-centred Values, Fairness, Privacy and Security, Reliability and Safety, Transparency and Explainability, Contestability, and Accountability.
Each principle has one principle question, which can be used to get high-level insights and understanding of RAI practices and risks associated with the principle. The principle question includes a set of sub-questions which are distributed across risk categories within the principle.
Risk categories were identified by conducting concept mapping to effectively comprehend relevant risk themes and concepts [24]. The concept mapping approach was used to analyze and group related risk questions, thereby identifying AI risk themes. Through this process, we identified gaps and areas for improvement (e.g., missing information, duplicated categories, structural issues), clarified the relationships between different risk themes, and provided a comprehensive overview of the key areas of concern for AI systems. Each category is exclusively associated with one principle. As shown in Figure 2, principle 1 has been further divided into multiple categories, each representing a specific aspect of the principle. The categories act as common risk themes that are clear, specific, and relevant to the AI ethics principle they represent. Additionally, a category contains one or more sub-categories, providing even more granularity and specificity to the questions and facilitating a better understanding of the different dimensions of each category.
Figure 3 represents the hierarchical organization of 26 risk themes and their distribution across eight AI principles. The concept map aims to visualize the structure of the RAI Question Bank in a clear and logical manner, promoting a comprehensive understanding of the question bank. Detailed concept maps for each principle, including the relationships between concepts, are presented in the following sections.
Figure 3: The risk concept map for RAI; It comprises 26 risk themes identified from seven AI frameworks.
@@
Subsequent questions are organized into second and third levels. Level 2 questions are typically more specific than level 1 questions and provide additional detail and context. These questions target managers and individuals (e.g., team members) with specific roles in overseeing AI-related projects. Level 3 questions are even more detailed and enables users to further break down the key risks. These questions are aimed at practitioners who require a detailed understanding of the technical and operational aspects of AI risk assessment.
The following shows the example questions at different levels.
@@
Figure 4: The distribution of risk questions and the source frameworks across the eight AI ethics principles.
@@
The source frameworks, including the EU AI Act and ISO standards, contribute to the distribution of the RAI Question Bank risk questions across the eight AI ethics principles. As depicted in Figure 4, while most frameworks cover five or six principles, the ISO standard particularly concentrates on Accountability as it primarily addresses management aspects.
In the following sections, we will describe in detail how the RAI Question Bank is organized by the eight AI ethics principles. This includes key objectives, risk themes, a concept map showing the relationships between these themes, and example questions for each principle.

### v2

@@ 1 Introduction
A recent report found that while many companies view AI as a promising technology and actively pursue AI opportunities, only 10% of those surveyed have publicly announced their Responsible AI (RAI) policies [2]. This highlights that many companies are still concerned about their RAI maturity level. Moreover, there have been numerous incidents involving AI [3]. AI incidents are not limited to specific industries like finance and healthcare. They have emerged across all sectors, raising various concerns in areas such as privacy, bias, and safety.
@@
There are a number of industrial RAI frameworks to support the implementation of RAI principles. In the US, the NIST AI Risk Management Framework [6] provides four functions such as MAP, GOVERN, MEASURE and MANAGE to manage AI risks. Recently, NIST has also released a new draft framework for Generative AI (GenAI) [7] , which encompasses 12 potential risks of GenAI and provides comprehensive RAI practices to address those risks. In Europe, the EU Assessment List for Trust AI Framework [8] has been used as AI ethics guidelines, which aligns with seven RAI principles (e.g., privacy, transparency, accountability) . In academia, researchers have increasingly focused on RAI practices and risk assessment for AI.
@@
While existing frameworks provide valuable guidance, they often remain high-level and struggle to bridge the gap between policy goals and practical implementation in AI development. Moreover, there is a lack of a comprehensive AI risk assessment framework that integrates interconnected, layered risk questions and aligns with other standards and frameworks [5]. Existing RAI frameworks often focus on specific aspects, failing to address the holistic nature of RAI principles. Liao et al. [9, 10] developed a question bank specifically focused on transparency and explainability, overlooking other crucial principles. Similarly, the US Department of Energy’s AI Risk Management Playbook [11] provides extensive guidance on identifying AI risks and recommending mitigations, but it lacks integration with diverse standards, frameworks, and related resources.
@@
To address current challenges, we present a novel RAI question bank developed based on a systematic literature review and in-depth survey on existing RAI frameworks. The initial results in 2023 were suitable for general and broad applications [12]. A key advancement of this study is the integration of a systematic approach that links lower-level risk questions to higher-level ones and related themes to prevent siloed and fragmented assessments, ensuring a cohesive evaluation process across different levels of the organization. Additionally, it is to accommodate different levels of stakeholders (e.g., executives, senior managers, developers) and is mapped to the system life-cycle stages. This allows users to focus on their specific contexts and interests while maintaining a comprehensive system perspective, thus enhancing both the relevance and effectiveness of the risk assessments.
@@
Although the RAI Question Bank relies primarily on yes/no questions, which risk fostering a checkbox mentality (i.e., the presence of low-quality artifacts or processes) if not carefully managed, we emphasize the importance of incorporating quality metrics and evidence-based responses. This means that the questions need to have quality aspects and answers are not just yes/no but require evidence, and lower-level answers collectively provide a crude metric for higher-level questions. Quality metrics and measurement are not the primary focus of this study.Yet, we have examined these considerations through a case study (Section 5.2) where we tested and refined the approach to ensure more meaningful and reliable assessments.
@@
The paper is organized as follows. Section 2 provides background information. Section 3 describes the overall design and development approach of the RAI Question Bank. We then present the Responsible AI Question Bank, offering valuable insights and serving as a fundamental resource for AI risk assessment (Section 4). Section 5 includes two case studies to illustrate its practical application for real-world use. Section 6 covers the practical implications of this study, including how the question bank can be applied across various purposes and contexts, such as compliance checks and risk assessments of AI agents like Auto-GPT [13]. We conclude this study in Section 7.
2 Background
@@
  • NIST AI Risk Management Framework [6]: The NIST AI RMF offers a flexible and voluntary framework applicable across various sectors and use cases. It enhances AI system trustworthiness by focusing on four main functions: Govern, Map, Measure, and Manage. Developed through an inclusive and transparent process, the framework supports existing AI risk management initiatives and adapts to the evolving AI landscape, enabling organizations to implement it to different extents for societal benefit while minimizing risks.
@@
  • Microsoft Responsible AI Impact Assessment Template and Guide [15, 16]: Microsoft’s resources support its Responsible AI Standard [17]. They enable AI development teams to explore and understand the impacts of their AI systems on stakeholders, intended benefits, and potential harms from the earliest design stages.
@@
  • EU Assessment List for Trustworthy Artificial Intelligence [8]: ALTAI helps organizations self-assess the trustworthiness of their AI systems. This checklist translates key requirements from the Ethics Guidelines for Trustworthy AI [18] into actionable steps, ensuring alignment with principles to mitigate risks and promote responsible AI development and deployment.
@@
  • Canada Algorithmic Impact Assessment [19]: This tool helps organizations evaluate the potential impacts of AI systems, especially in public services. It includes a detailed questionnaire to identify and mitigate risks associated with automated decision systems, focusing on bias, privacy, and transparency, and ensures compliance with ethical and legal standards.
@@
  • Australia NSW AI Assurance Framework [20]: This framework ensures the safe, ethical, and responsible use of AI technologies in government services. It includes AI governance, risk management, and mitigation strategies, emphasizing stakeholder engagement and continuous monitoring to maintain compliance with ethical standards. It also addresses emerging AI risks, such as those posed by generative AI, ensuring AI projects meet community expectations and uphold public trust.
AI risk assessments should account for the evolving landscape of AI regulations and standards such as the EU AI Act [21] and ISO/IEC 42001:2023 [22].
@@
The EU AI Act establishes a comprehensive regulatory framework for AI within the European Union. It classifies AI systems based on their risk levels: unacceptable, high, limited, and minimal risk. Unacceptable risk AI applications are banned, while high-risk applications must meet stringent requirements for transparency, quality, and security, and undergo conformity assessments. Additionally, it establishes the European Artificial Intelligence Board to ensure compliance and promote cooperation among member states. ISO/IEC 42001:2023 specifies requirements for establishing, implementing, maintaining, and continually improving an AI management system within organizations. It is designed for entities providing or utilizing AI-based products or services, ensuring responsible development and use of AI. The standard focuses on quality management and reliability, guiding organizations in managing the issues and risks associated with AI technologies. This systematic approach facilitates consistency with other management system standards related to quality, safety, security, and privacy.
@@
Recent efforts to improve the safety of advanced AI include several notable works. Partnership on AI released guidance for the safe deployment of foundation models, targeting model developers [30]. The AI Verify Foundation and Singapore’s Infocomm Media Development Authority published the Model AI Governance Framework for Generative AI, providing high-level guidance on the design, development, and deployment of generative AI models [31]. On April 29, 2024, NIST released a draft of the AI RMF Generative AI Profile, guiding organizations in managing risks unique to generative AI [7]. The EU AI Act also covers general-purpose AI like generative models, imposing specific transparency and evaluation requirements. The UK AI Safety Institute has published several resources, including Inspect [32], a platform for evaluating LLMs, and is actively developing infrastructure to assess the safety and societal impacts of advanced AI. Recently, the AU NSW Government updated its Artificial Intelligence Assessment Framework [20]. The revised framework, formerly known as the NSW Artificial Intelligence Assurance Framework, now categorizes generative AI-based solutions as "elevated risk," requiring additional and stricter assessments.
@@ 3 Methodology
This study was conducted in five dedicated phases from 2022 to 2024 (Figure 1). First, a systematic mapping study was performed to understand the state-of-the-art in AI risk assessment and select reference frameworks [5]. We selected five frameworks from the results, scrutinized, and extracted all included AI risk questions. The collected questions were synthesized and used to develop a comprehensive and holistic question bank for AI risk assessment [12]. We then conducted a case study with eight AI development projects. Some selected questions were tailored for each project and used for interviews with team members. We gained insights and feedback from the participating project teams. Based on the feedback, we improved the initial version of the question bank. As new AI regulations and standards are widely adopted, we incorporated the EU AI Act and ISO/IEC AI management standard to support organizations in achieving compliance. Finally, we evaluated the feasibility and applicability of the question bank through a case study of the ESG-AI framework development project.
Figure 1: Research methodology overview.
@@ 3.2 Question Bank Development
Based on the previous phase, five frameworks out of the 16 frameworks was chosen for further analysis and synthesis. The frameworks were selected according to i) global recognition, ii) inclusion of risk assessment questions, and, iii) representation of different regions and industry leaders. The selected frameworks include: US NIST AI Risk Management Framework [6], EU Trustworthy AI Assessment List [18], Canada Algorithmic Impact Assessment [19], Australia NSW AI Assurance Framework [20], and Microsoft Responsible AI Impact Assessment Guide [16]. By investigating these frameworks, we can comprehend and synthesize risk assessment solutions throughout various geopolitical contexts.
We began by extracting 382 questions for risk assessment, which were subsequently refined through an iterative process to consolidate questions on similar topics. This refinement resulted in a condensed question bank of 293 questions. We utilized Australia’s AI ethics principles [37] to categorize these questions. We structured the hierarchy of our question bank using a decision tree [38], organizing questions based on their levels (level 1-3) and sequences to ensure a coherent and comprehensive framework. To address potential gaps in technical/practical detail at lower level questions, we recognized that some lower-level (level 2 and 3) questions can be inferred from broader high-level frameworks. We made inferences to generate or extract these detailed questions where necessary. Concept mapping [39] was employed to identify common RAI themes, allowing us to further analyze and group the questions.
@@ 3.3 Case Study
The initial case study (Phase 3) with eight AI projects was conducted in 2023. We analyzed these projects to understand their contexts and selected 3 to 5 questions from our question bank for each AI ethics principle to prepare interviews. In parallel, we prepared a dedicated risk register template, including risk ID, category, title, description, causes, and selected interview questions derived from the question bank. For each project, we scheduled a 1.5-hour interview. We synthesized and analyzed the interview notes to identify and assess the AI ethics risks the project teams might encounter. We also asked the interviewees for feedback on the question bank for further refinement. We conducted two rounds of interviews in 2023 and consolidated the feedback to improve our question bank. The second case study (Phase 5) was conducted with the ESG-AI investor framework development project, following the question bank improvement (Phase 3), based on the initial case study results and changes in the external AI regulation landscape. This case study engaged 28 companies using and developing AI. Considering the companies’ characteristics and contexts, we selected and customized 5-10 questions for the interviews. In the ESG-AI framework development, 42 questions from the RAI Question Bank were selected to contribute to a deep dive assessment of companies. During the project, we received user feedback, improved the RAI Question Bank accordingly, and discussed the next steps.
@@ 4.1 Overview
This section presents the RAI Question Bank which is a set of risk questions designed to support the ethical development of AI systems. It addresses the needs of various users, including C-level executives, senior managers, developers, and others.
@@
Figure 2 presents the overall architecture of the RAI Question Bank. It is based on AI ethics principles which serve as an overarching category for the question bank. Specifically, we have selected Australia’s AI ethics principles [37] to comprehensively cover responsible AI concerns. The principles closely align with similar principles [40, 41] from around the world. The principles comprise eight principles such as Human, Societal and Environmental Wellbeing, Human-centred Values, Fairness, Privacy and Security, Reliability and Safety, Transparency and Explainability, Contestability, and Accountability. Each principle has one principle question, which can be used to get high-level insights and understanding of RAI practices and risks associated with the principle. The principle question includes a set of sub-questions which are distributed across risk categories within the principle.
@@
Categories were identified by conducting concept mapping to effectively comprehend relevant RAI concepts [42]. The concept mapping approach was used to analyze and group related risk questions, thereby identifying RAI themes. Through this process, we identified gaps and areas for improvement (e.g., missing information, duplicated categories, structural issues), clarified the relationships between different themes, and provided a comprehensive overview of the key areas of concern for AI systems. Each category is exclusively associated with one principle. As shown in Figure 2, principle 1 has been further divided into multiple categories, each representing a specific aspect of the principle. The categories act as common RAI themes that are clear, specific, and relevant to the AI ethics principle they represent. Additionally, a category contains one or more sub-categories, providing even more granularity and specificity to the questions and facilitating a better understanding of the different dimensions of each category.
Figure 3 represents the hierarchical organization of RAI risk categories including 26 categories and 65 sub-categories and their distribution across eight AI principles. The figure aims to visualize the structure of the RAI Question Bank in a clear and logical manner, providing a comprehensive overview of the taxonomy of RAI risk assessment areas derived from seven existing AI frameworks. This hierarchical representation not only facilitates the understanding of the relationships between AI ethics principles, categories, and sub-categories but also enables a systematic exploration of the key areas within RAI risk assessment. By organizing these elements visually, the figure serves as a guide for navigating complex ethical considerations, making it easier to identify how different risk areas are interconnected and aligned with existing AI frameworks. Additionally, it includes links between each sub-category and its corresponding source frameworks, illustrating how the seven existing AI frameworks have contributed to and informed the development of this taxonomy. As depicted in the figure, while most frameworks provide comprehensive coverage of the categories, the ISO standard emphasizes Accountability, focusing primarily on management aspects. Similarly, the EU AI Act places greater emphasis on Accountability and Transparency. This visual representation allows for a clear understanding of the provenance of each sub-category, demonstrating the alignment and integration of various frameworks in shaping the RAI Question Bank. Detailed concept maps for each principle, including the relationships between concepts, are presented in the following sections.
Figure 3: The RAI Question Bank categorization, comprising 26 categories and 65 sub-categories, identified from seven AI frameworks.
@@
Subsequent questions are organized into second and third levels. Level 2 questions are typically more specific than level 1 questions and provide additional detail and context. These questions target managers and individuals (e.g., team members) with specific roles in overseeing AI-related projects. Level 3 questions are even more detailed and enables users to further break down the key risks. These questions are aimed at practitioners who require a detailed understanding of the technical and operational aspects of AI risk assessment. The following shows the example questions at different levels.
@@
The following sections describe in detail how the RAI Question Bank is organized by the eight AI ethics principles.